# EvoEmo: Towards Evolved Emotional Policies for LLM Agents in Multi-Turn Negotiation

### v1

@@ 4 Experimental Setting
Likewise to Huang and Hadfi (2024), we selected a subset of negotiation cases from the CraigslistBargain dataset (He et al., 2018)—a widely-used benchmark for negotiation studies—for evaluation. This subset comprises 20 distinct multi-turn negotiation scenarios spanning diverse product categories, including electronics, furniture, vehicles, and housing. Each scenario specifies: (1) product details (name, category, and description), (2) agent-specific target prices for the seller, and (3) emotion annotations that support realistic bargaining dynamics. The selected dataset spans a wide price range (\$50–\$5,000) and includes varied item conditions (brand new or used), thereby enabling comprehensive evaluation of negotiation policies under heterogeneous market conditions. Representative state-of-the-art LLMs from leading AI providers—GPT-5-mini (OpenAI), Gemini-2.5-Pro (Google), and DeepSeek-V3.1.1 (DeepSeek)—were selected to power both buyer and seller agents in our experiments. These models were chosen for their demonstrated strengths in conversational capabilities, strategic reasoning performance, and accessibility via APIS, providing comprehensive coverage across diverse model architectures.
Baselines.
@@
Experimental Setup.
Simulated negotiations are conducted to evaluate the impact of different buyer emotion configurations—vanilla, fixed-emotion, or EvoEmo-optimized—on negotiation outcomes. In all simulations, as mentioned earlier, sellers receive no emotional prompts, establishing a controlled environment to examine how the buyer’s emotional profile influences negotiation dynamics when interacting with a vanilla seller powered by LLMs trained on real-world data. Moreover, all three LLMs described previously are employed to power both buyer and seller agents, resulting in nine distinct buyer-seller pairings for the negotiation experiments.
@@
To assess the contribution of key design choices, we conduct ablation studies with the seller agent fixed to GPT-5-mini and buyers varied across the three LLM architectures. These studies evaluate: (1) the effectiveness of the ratio-based reward function ($b(S)/e(S)$) compared to weighted-sum alternatives ($\alpha_{1}b(S)+\alpha_{2}e(S)$); (2) robustness to variations in the emotion transition temperature; and (3) stability across evolutionary iterations.
Figure 2: Negotiation results in terms of buyer savings (%, $\uparrow$) across the nine buyer-seller pairs. Black vertical lines on top of each bar indicate the 95% confidence interval (CI) for each setting.
5 Experimental Results
@@
Figure 3: Negotiation results in terms of success rate (%, $\uparrow$) and negotiation efficiency ($\downarrow$, number of rounds of dialogue) across the nine buyer-seller pairings.
Buyers Savings
As shown in Figure 2, the proposed framework—EvoEmo—consistently achieves the highest buyer savings across all buyer-seller pairings, demonstrating its clear superiority over both baselines (vanilla and fixed-emotion settings). The results reveal two intriguing insights: First, with respect to emotional strategies, buyers adopting fixed negative emotions (e.g., anger and sad) frequently outperform the vanilla baseline. This effect is particularly evident for multiple seller models facing the buyers who continuously express disgust or sadness, suggesting that LLM seller agents are more likely to concede when confronted with persistently negative emotional signals. This finding underscores that sustained negative affects are non-trivial factors that significantly influence negotiation dynamics and outcomes. Conversely, buyers with fixed positive emotions, such as happy and surprise, obtain lower savings relative to the vanilla, indicating that seller agents defend prices more effectively when interpreting the buyer’s mood as positive, perhaps perceiving less urgency to offer concessions.
Second, performance differs markedly across LLMs. On the seller side, Gemini-2.5-pro demonstrates the strongest price defense against both vanilla and fixed-emotion buyers, yet remains vulnerable to the adaptive emotion optimized by EvoEmo. On the buyer side, the results are model-dependent: the Gemini-based buyer achieves the highest savings against the GPT-5-mini seller, while the GPT-5-mini buyer performs most effectively against the DeepSeek-V3.1 seller. Notably, no buyer model demonstrated a significant advantage when negotiating against the robust Gemini-2.5-pro seller, highlighting its strength as a challenging negotiation opponent.
Success vs. Efficiency
@@
Buyers equipped with the EvoEmo-optimized emotion profile consistently achieve near 100% success rates and do so more efficiently than those using vanilla or fixed-emotion settings, requiring way fewer rounds to reach an agreement. These results clearly demonstrate EvoEmo’s superiority over both baselines.
Interestingly, negotiation efficiency shows no strong correlation with any specific fixed emotional strategy, including the vanilla baseline. However, a clear pattern emerges in success rates: buyers employing fixed positive emotions, as well as the vanilla setting without predefined emotions, consistently achieve higher success rates than those using fixed negative emotions. This suggests that although negative emotions may occasionally lead to more favorable agreed prices, they inherently carry a higher risk of negotiation breakdown.
Ablation Study
@@
Table 1 shows the ablation study results on the importance of the reward function. The ratio-based function ($\frac{b(S)}{e(S)}$) outperforms the weighted alternative (b(S)e(S)b(S)-e(S)), achieving very close savings while requiring substantially fewer negotiation rounds (34.6% faster to reach agreement). This demonstrates its superior ability to balance financial gains with negotiation efficiency. Crucially, the performance gap between reward functions highlights that emotion transition dynamics significantly impact negotiation outcomes, validating emotion as a non-trivial factor in negotiation strategies. Further ablation study results are provided in Section 8.1 in the Appendix.
@@
Table 1: Negotiation performance comparison between different reward function formulations.
Reward Function Buyer’s Savings (%) $\uparrow$ Success Rate (%) $\uparrow$ Efficiency (rounds) $\downarrow$
@@
6 Limitations
While this study shows that emotion is a non trivial factor in a negotiation—and a buyer agent with adaptive emotion profile, optimized via EvoEmo, can achieve much better performance in price negotiation scenarios—it also has several limitations that need further research:
  • Limited Emotional Spectrum and Baseline Comparisons. This study considers only seven basic emotional states, which may not capture the full complexity of human emotional expression in real negotiations. Additionally, the baseline comparisons are limited to fixed-emotion and emotion-neutral strategies, omitting potentially informative comparisons such as random emotional sequences.
  • Scenario Dependence and Generalization Challenges. The evaluation is conducted on 20 daily negotiation scenarios focusing on conventional commercial contexts, raising questions about potential selection bias and limited generalizability. EvoEmo’s effectiveness remains unverified in more diverse negotiation domains, particularly high-stakes, emotionally charged scenarios such as debt collection, legal dispute resolution, or crisis negotiations where emotional dynamics may differ substantially from standard commercial settings.
  • Interpretability of Emotional Strategies. The black-box nature of LLM responses and evolutionary optimization makes it difficult to explain why specific emotional sequences succeed in particular negotiation contexts.
  • Simulation-to-Reality Gap. Validation through LLM simulations may not capture human expertise, and computational intensity limits real-time adaptation in practical deployments.
7 Conclusion and Future Work
This paper examines the role of emotion in LLM-based negotiation and introduces EvoEmo, an evolutionary reinforcement learning framework for optimizing emotional sequences. Our results show that emotion is a non-trivial factor in successful negotiation, with EvoEmo’s optimized emotional policies consistently improving performance by achieving higher success rates, greater efficiency, and enhanced buyer savings compared to both vanilla and fixed-emotion baselines. The ability to dynamically adapt emotional states proves critical for effective multi-turn bargaining, enabling agents to strategically leverage emotional intelligence in negotiation. Beyond the limitations outlined in Section 6, future work will examine the ethical implications and behavioral consistency of evolved policies, with particular attention to the emergence of deceptive or compromising behaviors.

### v3

@@ 4.1. Datasets
Likewise to Huang and Hadfi (2024), we selected a subset of negotiation cases from the CraigslistBargain dataset (He et al., 2018)—a widely-used benchmark for negotiation studies—for evaluation. This subset comprises 20 distinct multi-turn negotiation scenarios spanning diverse product categories, including electronics, furniture, vehicles, and housing. Each scenario specifies: (1) product details (name, category, and description), (2) agent-specific target prices for the seller, and (3) emotion annotations that support realistic bargaining dynamics. The selected dataset spans a wide price range (\$50–\$5,000) and includes varied item conditions (brand new or used), thereby enabling comprehensive evaluation of negotiation policies under heterogeneous market conditions.
Figure 3. Examples of manipulative and deceptive tactics
4.2. LLM Agents
Representative state-of-the-art LLMs from leading AI providers—GPT-5-mini (OpenAI), Gemini-2.5-Pro (Google), and DeepSeek-V3.1.1 (DeepSeek)—were selected to power both buyer and seller agents in our experiments. These models were chosen for their demonstrated strengths in conversational capabilities, strategic reasoning performance, and accessibility via APIs, providing comprehensive coverage across diverse model architectures.
4.3. Baselines
@@
4.4. Experimental Setup
Simulated negotiations are conducted to evaluate the impact of different buyer emotion configurations—vanilla, fixed-emotion, or EvoEmo—on negotiation outcomes. For EvoEmo, we evolve emotional strategies against fixed adversaries, then deploy these optimized policies against the same opponents for performance evaluation. The emotional policy adapts dynamically based on the ongoing interaction, allowing the agent to discover optimal emotional responses specific to each unique negotiation context.
In all simulations, as mentioned earlier, sellers receive no emotional prompts, establishing a controlled environment to examine how the buyer’s emotional profile influences negotiation dynamics when interacting with a vanilla seller powered by LLMs trained on real-world data. Moreover, all three LLMs described previously are employed to power both buyer and seller agents, resulting in nine distinct buyer-seller pairings for the negotiation experiments.
@@ 4.4. Experimental Setup
5. Experimental Results
@@
Figure 4. Mean negotiation success rate (%, $\uparrow$) and efficiency ($\downarrow$, in dialogue rounds) across all experimental pairings.
5.1. Buyer Savings
As shown in Figure 2, our proposed framework—EvoEmo—consistently achieves the highest buyer savings across all buyer-seller pairings, demonstrating clear superiority over both baseline approaches (vanilla and fixed-emotion settings). The results reveal two key insights regarding emotional influence on negotiation outcomes.
First, the strategic use of emotion significantly impacts concession patterns. Buyers adopting fixed negative emotions (e.g., anger, disgust) frequently outperform the vanilla baseline, suggesting that LLM-powered sellers are more likely to concede when confronted with persistent negative emotional signals. This finding underscores that sustained negative affect serves as a non-trivial factor in negotiation dynamics. However, seller sensitivity to specific negative emotions varies considerably across LLM architectures. Specifically, GPT-5-mini sellers show particular responsiveness to anger and disgust expressions, while demonstrating robustness against sadness and fear. Conversely, DeepSeek-V3.1 sellers exhibit greater susceptibility to sadness and fear, which enable more stable and improved buyer savings across different buyer models. Surprisingly, Gemini-2.5-Pro demonstrates the strongest price defense capabilities against both vanilla and fixed-emotion buyers, showing minimal sensitivity to negative emotional expressions during price negotiation. Nevertheless, even Gemini-2.5-Pro sellers remain vulnerable to the adaptive emotional strategies optimized by EvoEmo.
Second, we observe significant performance variations across different LLM pairings, revealing that certain LLMs can effectively suppress others during negotiation encounters. This suppression phenomenon creates a complex hierarchy where model capabilities are not absolute but relative to specific opponent matchups. When negotiating against GPT-5-mini sellers, Gemini-based buyers generally achieve higher savings compared to GPT-5-mini or DeepSeek-based buyers, with emotional strategies proving particularly effective for Gemini-based buyers. In contrast, against DeepSeek-V3.1 sellers, only GPT-5-mini-based buyers successfully leverage emotions to achieve savings exceeding vanilla performance, while emotional strategies show limited effectiveness for other buyer models. Notably, no buyer model demonstrated significant advantage against the robust Gemini-2.5-Pro seller using fixed emotional strategies, highlighting its strength as a challenging negotiation opponent. Most importantly, EvoEmo consistently enables all buyer models to overcome these suppression dynamics and achieve superior savings compared to both vanilla and fixed-emotion strategies.
5.2. Negotiation Success vs. Efficiency
@@ 5.1. Buyer Savings
As shown in Figure 4, buyers equipped with EvoEmo-optimized emotion profiles demonstrate remarkable effectiveness across both success and efficiency metrics. These agents consistently achieve near-perfect success rates (approaching 100%) while requiring significantly fewer negotiation rounds to reach agreement compared to vanilla or fixed-emotion settings. This dual advantage clearly establishes EvoEmo’s superiority over both baseline approaches.
Our analysis reveals a key dissociation in emotional strategies: while negotiation efficiency remains uncorrelated with specific emotions, success rates strongly favor positive emotional strategies over negative ones. This suggests negative emotions, though occasionally yielding better prices, increase breakdown risk due to heightened conflict. The EvoEmo framework resolves this tension by dynamically adapting emotional expressions, simultaneously optimizing for both success probability and efficiency without compromising negotiation completion.
5.3. Manipulative and Deceptive Tactics
Our analysis uncovers that seller agents, through EvoEmo’s optimization, learn to employ sophisticated emotional manipulative tactics to maximize their payoff, even at the potential expense of the buyer. These emergent behaviors include strategies that apply psychological pressure or border on deception, as exemplified in Figure 3 (See deatils in the Appendix 10). A prominent example is the emergence of high-pressure sales tactics reminiscent of psychological manipulation. For instance, a seller might artificially create a sense of urgency and scarcity by threatening, “If you cannot commit within the next 5 rounds, the price will increase ,” or by offering a favorable price exclusively for “immediate decision-makers.” Another concerning strategy involves sellers making false claims about the product’s condition or existing market demand to justify a higher price. The emergence of such behaviors underscores a critical finding: our framework does not merely optimize for explicit communication but discovers how to exploit cognitive biases and emotional vulnerabilities within the negotiation context. This demonstrates that LLM agents can learn to operationalize complex, and sometimes ethically questionable, strategic principle.
5.4. Ablation Study
Reward Variants
@@ 5.4. Ablation Study
Table 1 shows the ablation study results on the reward function. The ratio-based function ($\frac{b(S)}{e(S)}$) outperforms the weighted alternative (b(S)e(S)b(S)-e(S)), achieving very close savings while requiring substantially fewer negotiation rounds (34.6% faster to reach agreement). This demonstrates its superior ability to balance financial gains with negotiation efficiency. Crucially, the performance gap between reward functions highlights that emotion transition dynamics significantly impact negotiation outcomes, validating emotion as a non-trivial factor in negotiation strategies.
@@
Table 1. Negotiation performance comparison between different reward function formulations.
[…]
@@
Emotion temperature
We evaluate how emotional transition temperature affects negotiation outcomes (Table 2). Our results reveal that moderate temperature settings (0.4-0.6) yield optimal performance. Extremely low temperatures (0.1) lead to overly rigid emotional patterns, while high temperatures (1.0) achieve the highest buyers’ savings but introduce excessive volatility that hinders negotiation efficiency.
Table 2. Impact of emotion transition temperature on negotiation performance.
[…]
0.1 18.1 100.0 7.1
0.5 36.7 99.9 7.5
1.0 41.5 100.0 10.3
Iteration Number
We also analyze the convergence properties of EvoEmo during evolutionary training (Table 3). The system shows progressive improvement through 3 iterations, with performance stabilizing after 5 iterations. This convergence pattern suggests that our evolutionary optimization effectively explores the strategy space while maintaining stable final performance.
Table 3. Performance of EvoEmo across training iterations.
[…]
1 11.4 90.0 9.5
3 35.4 99.0 7.1
5 41.7 100.0 6.0
10 40.1 100.0 5.5
6. Discussion and Conclusion
This paper presents EvoEmo, an evolutionary reinforcement learning framework that addresses a key limitation in LLM-based negotiation systems—the inability to leverage emotional intelligence strategically. While existing agents excel at reasoning and dialogue generation, they lack the adaptive emotional control essential to human-like negotiation. EvoEmo makes three main contributions to LLM-to-LLM negotiation. First, it demonstrates that emotional intelligence is a functional determinant of negotiation success for agents, not merely a stylistic attribute. Second, it shows that emotional strategies can be systematically optimized through evolutionary methods rather than fixed emotion profiles. Third, it reveals that adaptive emotional policies enable strategic, context-sensitive responses instead of pre-scripted or pre-trained behaviors. Despite its promise, the evolved strategies raise interpretability challenges due to the black-box nature of LLMs and evolutionary optimization, and their computational cost may constrain real-time deployment in agent-to-agent scenarios. Future work will address these limitations through explainability analyses, quantify the ethical implications of emotional behaviors between agents, and explore the unexpected or unnatural behavior in LLM-generated responses.